. An Engineering Company is considering an investment proposal to install new milling controls. The project cost is `. 50,000. The facility has life of 5 years and no salvage value. The company’s tax rate is 55 %. The estimated cash flows before tax (CFBT) from the proposed investment proposal are as follows:                                                                                4

Year​​​​​​CFBT (`)

1​​​​​​10,000

2​​​​​​11,000

3​​​​​​14,000

4​​​​​​15,000

5​​​​​​25,000

Compute the following:

i) Pay- Back Period

ii) Average rate of return

iii) Net present value at 10% discount rate

iv) Profitability index at 10% discount rate



1
Expert's answer
2021-02-17T17:21:52-0500

Initial Cost of Investment = 50000

Yearly Depreciation = 50000/5 = 10000

Since Profit before Tax is given in the question, it is assumed that depreciation has already been deducted , therefore depreciation is to be added back for calculating cash flows



Payback Period - 3 + 4250/16750 = 3.25 Years

NPV = Sum of discounted Cash flows - initial investment

= 62407- 50000 = 12407

Average Rate of return can be calculated as follows = Sum of net Cash flows after tax = 83750

Net benefit in 5 years = 83750 - 50000 = 33750

Yearly benefit = 6750

So, ARR = 6750/50000 = 13.50%